40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Nov 18 22:49:06 2020, gautam, Update, ASC, Some thoughts about AS WFS electronics 
    Reply  Sun Nov 22 21:09:37 2020, gautam, Update, ASC, Planned mods for WFS head aLIGO_wfs_v5_40m.pdfTFs.pdfnoise.pdf
       Reply  Wed Nov 25 18:18:41 2020, gautam, Update, ASC, Planned mods for WFS head anodeVsOutput.pngsensitivity.png
          Reply  Sat Nov 28 21:44:53 2020, rana, Update, ASC, Planned mods for WFS head 
          Reply  Thu Dec 10 22:46:57 2020, gautam, Update, ASC, WFS head assembled aLIGO_wfs_v5_40m.pdfTF_meas.pdf
             Reply  Thu Dec 17 15:23:56 2020, gautam, Update, ASC, WFS head characterization oNoise.pdfZ_Hi.pdfZ_Low.pdfIMG_9030.jpgiNoise.pdf
    Reply  Wed Nov 25 18:30:23 2020, gautam, Update, ASC, Some thoughts about AS WFS electronics 
       Reply  Wed Dec 2 23:07:19 2020, gautam, Update, ASC, Electrical LO signal for AS WFS RF_Frequency_Source.pdf
          Reply  Thu Dec 3 11:02:35 2020, rana, Update, Electronics, Electrical LO signal for AS WFS 
          Reply  Thu Dec 3 13:55:52 2020, Koji, Update, ASC, Electrical LO signal for AS WFS 
             Reply  Thu Dec 3 20:38:46 2020, gautam, Update, ASC, Electrical LO signal for AS WFS RF_Frequency_Source.pdfdemodPath.pdfEOMpath.pdf
                Reply  Thu Dec 3 21:14:38 2020, Koji, Update, ASC, Electrical LO signal for AS WFS 
                   Reply  Thu Dec 3 21:44:49 2020, gautam, Update, ASC, Electrical LO signal for AS WFS 
                      Reply  Fri Dec 4 03:02:40 2020, Koji, Update, ASC, Electrical LO signal for AS WFS 
                         Reply  Fri Dec 4 22:41:56 2020, gautam, Update, ASC, Freq Gen Box revamp 
                            Reply  Sat Dec 5 20:44:35 2020, gautam, Update, ASC, Freq Gen Box re-installed 6x
                               Reply  Mon Dec 7 12:38:51 2020, gautam, Update, IOO, IMC loop char OLTFcomparison.pdfdemodMons.pdfOLTFcomparison.pdf
                               Reply  Mon Dec 7 22:54:30 2020, gautam, Update, LSC, Modulation depth measurement modDepth.pdf
                                  Reply  Sat Jan 16 18:59:44 2021, gautam, Update, LSC, Modulation depth measurement PLL.pdfmodDepth.pdfmodProfile.pdfEOMpath_postMod.pdf
Message ID: 15769     Entry time: Sat Jan 16 18:59:44 2021     In reply to: 15715
Author: gautam 
Type: Update 
Category: LSC 
Subject: Modulation depth measurement 

I decided to analyze the data I took in December more carefully to see if there are any clues about the weird LSC sensing.

Attachment #1 shows the measurement setup.

  • The PSL shutter was closed. All feedback to both lasers was disconnected during the measurement. I also disabled the input switch to the FSS Box - so the two laser beams being interfered shouldn't have any modulations on them other than the free running NPRO noise and the main IFO modulations.
  • Everything is done in fiber as I had the beams already coupled into collimators and this avoided having to optimize any mode matching on the beat photodiode.
  • The pickoff of the PSL is from the collimator placed after the triply resonant EOM that was installed for the air BHD experiment.
  • The other beam is the EX laser beam, arriving at the PSL table via the 40m long fiber from the end (this is the usual beam used for ALS).
  • I didn't characterize precisely the PLL loop shape. But basically, I wasn't able to increase the SR560 gain any more without breaking the PLL lock. Past experience suggests that the UGF is ~20 kHz, and I was able to get several averages on the AG4395 without the lock being disturbed.

Attachment #2 shows the measured spectrum with the PSL and EX laser frequency offset locked via PLL.

  • The various peaks are identified.
  • There are several peaks which I cannot explain - any hypothesis for what these might be? Some kind of Sorensen pollution? They aren't any multiples of any of the standard RF sources. They are also rather prominent (and stationary during the measurement time, which I think rules out the cause being some leakage light from the EY beam, which I had also left connected to the BeatMouth during the measurement).
  • In the previous such characterization done by Koji, such spurious extra peaks aren't seen.
  • Also, I can't really explain why some multiples of the main modulation are missing (could also be that my peak finding missed the tiny peaks)?
  • The measuremet setup is very similar to what he had - important differences are 
    • Much of the optical path was fiber coupled.
    • Beat photodiode is NF1611, which is higher BW than the PDA10CF.
    • The second laser source was the Innolight EX NPRO as opposed to the Lightwave that was used.
    • The RF source has been modified, so relative phasing between 11 MHz and 55 MHz is different.

Fitting the measured sideband powers (up to n=7, taking the average of the measured upper and lower sideband powers to compute a least squares fit if both are measured, else just that of the one sideband measured) agains those expected from a model, I get the following best fit parameters:

\begin{align*} \Gamma_1 &= 0.193 \pm 0.004 \\ \Gamma_2 &= 0.246 \pm 0.008 \\ \phi &= 75.5^{\circ +17.5^{\circ}}_{\, -40.3^{\circ}} \end{align*}

To be explicit, the residual at each datapoint was calculated as

\Delta = \bigg| \frac{\rm{model}-\rm{measurement}}{\rm{model}}\bigg|^2.

The numbers compare favourably with what Koji reported I think - the modulation depths are slightly increased, consistent with the RF power out of the RF box being slightly increased after I removed various attenuators etc. Note the large uncertainty on the relative phase between the two modulations - I think this is because there are relatively few sidebands (one example is n=3) which has a functional dependence that informs on phi - most of the others do not directly give us any information about this parameter (since we are just measuring powers, not the actual phase of the electric field). 

Attachment #3 shows a plot of the measured modulation profile, along with the expected heights plugging the best fit parameters into the model. The size of the datapoint markers is illustrative only - the dependence on the model parameters is complicated and the full covariance would need to be taken into account to put error bars on those markers, which I didn't do. 

Attachment #4 shows a time domain measurement of the relative phasing between the 11 MHz and 55 MHz signals at the EOM drive outputs on the RF source box. I fit a model there and get a value for the relative phase that is totally inconsistent from what I get with this fit. 

Attachment 1: PLL.pdf  108 kB  | Show | Hide all | Show all
Attachment 2: modDepth.pdf  95 kB  Uploaded Mon Jan 18 18:51:56 2021  | Hide | Hide all | Show all
modDepth.pdf
Attachment 3: modProfile.pdf  151 kB  Uploaded Mon Jan 18 18:52:06 2021  | Show | Hide all | Show all
Attachment 4: EOMpath_postMod.pdf  183 kB  Uploaded Mon Jan 18 18:52:32 2021  | Show | Hide all | Show all
ELOG V3.1.3-